The specific heat capacity of a particular solid material at temperature close 0k is given by c=aT³. What is the S.I unit of the constant where T=temperature
1
Expert's answer
2020-01-24T05:17:27-0500

The SI unit of specific heat capacity is "Jkg^{-1}K^{-1}". The dimensions on both side of the expression "c = aT^3" must be equal. Therefore, the SI unit of constant "a" is:


"[a] = Jkg^{-1}K^{-4} = kgm^2s^{-2}kg^{-1}K^{-4} = m^2s^{-2}K^{-4}."

Answer:

"[a] = m^2s^{-2}K^{-4}."